25 workers are at a house project that makes them a month to finish in a flash what could possibly be the power of the workers if each worker works for 200N and then man power is increased by 60% to make the work efficient and effective at the same time, and the distance between the project site and the agency is 30 km from each other.


1
Expert's answer
2021-10-18T09:09:02-0400

Initial power of a worker who exerts a force of 200 N along 30 km during one month (30 days):


"P=\frac Wt=\frac{Fd}t=\frac{200·30000}{30·24·3600}=2.3\text{ W}."

Increased by 60%, it is


"P_i=1.6·2.3=3.7\text{ W}."